OpenStudy (anonymous):

Find the first, fourth, and eighth terms of the sequence. A(n)=-2*2^n-1

OpenStudy (jdoe0001):

hmmm \(\large \bf a_{\color{brown}{ n}}=2\cdot 2^{{\color{brown}{ n}}-1}\)

OpenStudy (jdoe0001):

\(\large \bf a_{\color{brown}{ n}}=2\cdot 2^{{\color{brown}{ n}}-1}?\) rather...is that... the formula

OpenStudy (jdoe0001):

\(\large { a_{\color{brown}{ n}}=2\cdot 2^{{\color{brown}{ n}}-1} \\ \quad \\ \begin{array}{llll} term&value \\\hline\\ a_{\color{brown}{ 1}}&a_{\color{brown}{ 1}}=2\cdot 2^{{\color{brown}{ 1}}-1}\to 2\cdot 2^0\to 2\cdot 1\to 2\\ a_{\color{brown}{ 4}}&a_{\color{brown}{ 4}}=2\cdot 2^{{\color{brown}{ 4}}-1}\to 2\cdot 2^3\\ a_{\color{brown}{ 8}}&a_{\color{brown}{ 8}}=2\cdot 2^{{\color{brown}{ 8}}-1}\to 2\cdot 2^7\\ \end{array} }\)

geerky42 (geerky42):

You can do this to make this problem easier: \(\Large 2\cdot2^{n-1}\rightarrow 2^1\cdot2^{n-1}\rightarrow 2^{n-1}+1=\boxed{2^n}\)

geerky42 (geerky42):

\(\Large 2^{n-1+1}\), not \(\Large 2^{n-1}+1\)

OpenStudy (anonymous):

Thanks.